### v1

@@ 1 INTRODUCTION
Future and current 21 cm intensity mapping telescopes such as CHIME [1], CHORD [2], HERA [3], HIRAX [4], LOFAR [5], MWA [6], SKA [7] [8] will measure the distribution of neutral hydrogen across a wide range of redshifts for a variety of science goals [9]. Isolating the cosmological 21 cm signal by removing bright radio foregrounds [10, 11, 12, 13, 14, 15] will require precision beam measurements to better than 1% uncertainty [16, 17, 18, 10, 19]. Calibration with well known radio point sources is difficult for stationary (un-steerable) drift scanning instruments, and one promising alternative is to use radio sources onboard drones as piloted calibration references [19, 20].
@@
CHIME is a radio interferometer operating between 400-800 MHz to measure redshifted neutral hydrogen. The cylindrical reflectors observe the sky by drift scanning with its large field of view. The resulting primary beam poses a unique challenge for drone beam mapping for several reasons: (1) the far-field is distant, at altitudes in excess of 1 km at all frequencies, (2) it is highly elliptical (1.2-2 East-West, and $\sim 120^{\circ}$ North-South) with frequency- and declination-dependent variations [30, 1].
In this paper we present measurements of the CHIME primary beam obtained with drone-based beam mapping techniques. We acquire these measurements at a declination comparable to existing solar [30] and holography [31] beam patterns. To detect the pulsing drone calibration signal we utilize the pulsar gating mode of CHIME to perform in-situ background subtraction. The resulting comparisons we present here provide independent validation of sky-based beam models and demonstrate the synchronization of the calibrator and receiver enabling future measurements of CHIME with a drone-based platform (e.g. polarization, repeatability, near-field to far-field).
Due to flight altitude restrictions, the flight took place squarely in the radiating near-field, hindering quantitative comparisons to the beams obtained from solar and holography datasets. Instead, we compare best-fit main beam parameters and other beam properties across frequency and declination to test the feasibility of these methods. These measurements will inform future near-field beam mapping efforts and near-field to far-field transformations of measured CHIME beams.
@@ 2.1 CHIME
To this end, CHIME is designed to observe the redshifted 21 cm emission of HI over a wide band between 400-800 MHz, corresponding to redshift $z=0.8-2.5$, considered a critical epoch for constraining competing models of dark energy vis-a-vis their impact on the expansion history. The CHIME instrument is described in more detail in [1], briefly, the telescope consists of four parabolic cylinders, 20 m wide (East-West) by 100 m long (North-South). Each cylinder is outfitted with 256 dual-polarization cloverleaf antennas separated by $\sim$30 cm within the inner 80 m length of the focal line. The telescope has no moving parts; the cylindrical reflectors provide a beam shape which is focused in the transverse East-West direction but extends nearly from horizon-to-horizon in the North-South direction. This accommodates a “driftscan” observing strategy; every 24 hours the telescope observes the entire sky accessible from the CHIME latitude due to the Earth’s rotation. This stationary design complicates beam calibration, as we cannot “point” the beam at a suitable calibrator. Alternatively, in this paper, we deploy a remotely controlled drone outfitted with a radiating payload as calibrator source to sample the instrument field of view.
@@ 2.3 CHIME Configuration for Drone Beammapping
The drone noise source is a single mode source seen by all CHIME feeds. After the $N^{2}$ correlation there is effectively an equal amount of information in each entry of the visibility matrix and to maximise the sensitivity of the measurements we would like to retain all this information. However this poses a severe data rate challenge. To avoid smearing within an integration frame we must reduce the accumulation time from the standard 10 s down to 2 s, but this increased data rate of around 50 TB/hour can neither be offloaded from the CHIME X-engine nor can any reasonable length of data be stored within the on-site storage. To get around this we used the existing capabilities of the CHIME Science Data Processor (SDP) system to achieve a significant compression of the data[1].
First we note that although there is signal in each element of the correlation matrix, for a single mode noise source all of this is captured within the first eigenmode of the matrix. Although any motion of the drone within an integration frame can generate additional modes, at 2 s cadence it is sufficient to record only a small number of the eigenvalues and vectors of the matrix to capture all of the information. Generating this eigendecomposition is computationally costly, and CHIME was designed to be able to do this once per 10 s frame. To fit within the computational footprint we only include data from a single cylinder of the instrument, which further reduces the output data rate. As we use an iterative algorithm for this calculation this gives only an $\sim N^{2}=16$ computational saving, rather than the $N^{3}$ increase we would expect for a full calculation.
The set of modified processing stages we perform on the individual X-engine nodes are:
  • Use the existing CHIME pulsar gating capability to take on-off differences of the noise source signal on the full $N^{2}$ visibility matrix. This is possible as both CHIME’s clock and that of the noise source are locked to the GPS PPS signal. This is then integrated to 2 s cadence.
@@
  • The existing eigendecomposition capability is used to calculate the largest 10 eigenvalues and their eigenvectors.
These eigenmodes as well as the full visibility matrix for cylinder C is then sent to the receiver node of the SDP system for the final processing, and writing of the following data products:
  • The visibility matrix is written out for a set of 64 frequencies to allow us to verify that the eigendecomposition extracted all of the information.
  • For the full set of frequencies we write out the eigenvalues and eigenvectors to allow full sensitivity analysis of the data, as well as measurements of phase variation.
@@
The gated power for each feed can be reconstructed using the strongest 10 eigenmodes from a single cylinder. The analysis in this paper uses the first eigenmode only, which sufficiently captures all of the drone signal while allowing for a lower noise floor than what is possible using the directly measured power contained in the autocorrelations.
@@ 2.4 Drone Flights
We performed ten flights over the CHIME telescope; in this paper we present results from the flight at the approximate declination of Virgo A, where independent beam measurements of CHIME from both holography data [31] and solar data [30] are available. Replicating the trajectory of Virgo A and the Sun through the CHIME primary beam required a 35 zenith angle drone flight relative to the array center. Due to altitude restrictions, we performed a low velocity flight at a conservative altitude to maximize the spatial resolution of the transit data. This flight was performed $\sim$220 m South from the center of the array, corresponding to distances between 180 - 250 m with respect to feed location along the focal line. Relative to the CHIME array center, the flight spanned -100 m to +100 m from East-to-West at an altitude of $\sim$307 m, in accordance with our flight ceiling limits from TransportCanada. The first flight was flown with a heading of $\sim 0^{\circ}$ [North] to measure co-polarizations corresponding to CHIME’s N (YY) polarization, and the second flight was flown with a heading of $\sim 90^{\circ}$ [East] to measure the orthogonal (XX) polarization. The typical distance from CHIME for these flights ranges from 350 m - 390 m. As a result, this flight plan corresponds to the near-field for all CHIME antennas (the CHIME far-field for a single feed is $>$1 km). The time for each polarization pass was 3.4 mins.
@@ 3.1 Map Projections
During data processing, each CHIME correlation frame is associated with a concurrent drone position measurement from the RTK unit, which expresses the drone location in latitude, longitude, and height coordinates. A conversion to a local Cartesian reference frame [^1] PyGeodesy is performed using the latitude, longitude, and height of a known reference point (in this instance, the center of the CHIME array provided by surveying).
CHIME beam parameters are typically presented in either celestial coordinates (Hour Angle and Declination) or orthographic coordinates (‘TelX’ and ‘TelY’) in which the coordinate TelX is always parallel to East and TelY to North, such that the beam is not distorted by sky curvature[30]. The linear array of the CHIME feeds along the cylinder results in each feed ‘seeing’ the drone at a different declination (see Figure 2). As a result, we compute the apparent position in RA/Dec or TelX/TelY coordinates using the drone location and relative position of each feed with respect to the center of the CHIME array. An example of one pass from the flight, in each of these coordinates, is shown in Figure 3
The mapping between telescope feed and declination crucially depends on the altitude of the drone flight. The altitude must be measured relative to the phase center of the array (where incident rays are focused onto the feedline). We model the phase center of each feed by extending the incident ray to an image of the feedline placed beneath the primary reflector at a distance of 5 m (the focal length). This defines the image plane. The height of the CHIME cylinders above ground is $\sim$1 m, thus the image plane sits $\sim$4 m below the ground surface. As a result, we expect to use an altitude of 307.6 m (corresponding to 311.6 m above the image plane), although we independently constrain the effective altitude in Section 3).
Figure 3: Coordinate conversions applied to data from a drone flight at 35 degrees of zenith angle. One transit across the beam of Cylinder C is shown true to scale in a local Cartesian coordinate system (left). The resulting transformations are shown for Cartesian (upper right), Celestial (middle right), and Orthographic (TelX/TelY) (lower right) for each feed using the color scale from Figure 2.
@@ 3.2 Available Datasets
Comparisons are possible between three independent data sets (a) Drone measurements; (b) Solar measurements; and (c) Holography measurements. The overlap between these data sources can be seen in TelX/TelY coordinates in Figure 4. In the drone flights described above, because each feed will map to a different declination, the flights mapped out declinations between 10 to 20 DEC and hour angles between -15 +15 HA. The drone data declination range overlaps with a portion of the data acquired during solar transits to map the CHIME beam shape, which range from -23 to +23 DEC. The drone data also overlaps with one source measured via holography, Virgo A at +12 DEC. All three data sets overlap at TelY $\sim-0.6$ which corresponds to a declination of $\sim 12^{\circ}$.
[…]
Figure 4: Beam characterization measurements have been calculated from three unique data sets acquired in different portions of CHIME’s field of view: Solar beam measurements are presented in a continuous semicircular slice. Holographic measurements of galactic radio sources appear as pale yellow arcs. Drone measurements are presented as grids with yellow outlines, while the drone measurement analyzed in this paper is outlined with a dashed yellow line. Comparison of the FWHM and Centroid measurements obtained from these data sets are shown in Figures 9 & 8.
The solar data is projected into the orthographic coordinate system (TelX, TelY) with primary axes aligned North-South and East-West. The advantage of this projection is that it orthogonalizes the features of the beam, and thus we use this projection throughout this paper. However, sources at constant declination such as the holographic measurements of Virgo A and the drone flight are curved in TelX/TelY (see Figure 4), which distorts sidelobe features outside $|\mathrm{TelX}|\gtrsim 0.25$. Further, because this range is far larger than the beamwidths, we found any correction to the centroid and FWHM fits of the primary beam are small compared to fitting errors.
Data Set Solar Holography Drone
@@
Declination () -23 – +23 (-12.1, 5.0, 12.5, +10 – +20
22.0, 40.75, 58.75)
Hour Angle () -60 – +60 -30 – +30 -15 – +15
Table 1: Survey coverage and parameter space of each data set. Virgo A for the holographic data set is in bold.

### v2

@@ 1 INTRODUCTION
Future and current 21 cm intensity mapping telescopes such as CHIME [1], CHORD [2], HERA [3], HIRAX [4, 5], LOFAR [6], MWA [7], SKA [8] [9] will measure the distribution of neutral hydrogen across a wide range of redshifts for a variety of science goals [10]. Isolating the cosmological 21 cm signal by removing bright radio foregrounds [11, 12, 13, 14, 15, 16] will require precision beam measurements to better than 1% uncertainty [17, 18, 19, 11, 20]. Calibration with well known radio point sources is difficult for stationary (un-steerable) drift scanning instruments, and one promising alternative is to use radio sources onboard drones as piloted calibration references [20, 21].
@@
CHIME is a radio interferometer operating between 400-800 MHz to measure redshifted neutral hydrogen (see [1] for more detail). The cylindrical reflectors observe the sky by drift scanning with its large field of view. The resulting primary beam poses a unique challenge for drone beam mapping for several reasons: (1) the far-field is distant, at altitudes in excess of 1 km at all frequencies, (2) it is highly elliptical (1.2-2 East-West, and $\sim 120^{\circ}$ North-South) with frequency- and declination-dependent variations [31, 1, 32].
In this paper we present measurements of the CHIME primary beam obtained with drone-based beam mapping techniques. We acquire these measurements in a region which overlaps with existing solar [31] and holography [33, 32] beam measurements. To detect the pulsing drone calibration signal we utilize the pulsar gating mode of CHIME to perform in-situ background subtraction. The resulting comparisons we present here provide independent validation of sky-based beam models and demonstrate the synchronization of the calibrator and receiver enabling future measurements of CHIME with a drone-based platform (e.g. polarization, repeatability, near-field to far-field).
Due to flight altitude restrictions, the flight took place squarely in the radiating near-field, hindering direct comparisons of the beam as a function of angle between drone data, solar data and holography data sets. Instead, we compare best-fit main beam parameters and other beam properties across frequency and declination to test the feasibility of these methods. These measurements will inform future near-field beam mapping efforts and near-field to far-field transformations of measured CHIME beams.
@@ 2.1 CHIME
To this end, CHIME is designed to observe the redshifted 21 cm emission of HI over a wide band between 400-800 MHz, corresponding to redshift z=0.82.5z=0.8-2.5, considered a critical epoch for constraining competing models of dark energy vis-a-vis their impact on the expansion history. The CHIME instrument is described in more detail in [1], briefly, the telescope consists of four parabolic cylinders, 20 m wide (East-West) by 100 m long (North-South). Each cylinder is outfitted with 256 dual-polarization cloverleaf antennas separated by $\sim$30 cm within the central 80 m length of the focal line. The telescope has no moving parts; the cylindrical reflectors provide a beam shape which is focused in the transverse East-West direction but extends nearly from horizon-to-horizon in the North-South direction. This accommodates a “driftscan” observing strategy; every 24 hours the telescope observes the entire sky accessible from the CHIME latitude due to the Earth’s rotation. This stationary design complicates beam calibration, as we cannot “point” the beam at a suitable calibrator. Alternatively, in this paper, we deploy a remotely controlled drone outfitted with a radiating payload as calibrator source to sample the instrument field of view.
@@ 2.2 Drone Instrumentation
The signal is transmitted from a single-polarization biconical antenna [^2] Aaronia Bicolog 30100, the same used in [21], with a frequency range operating between 30MHz-1GHz. A commercial bandpass filter was included in the analog chain to restrict the transmission signal to 400-800 MHz. The transmitting antenna is isolated from the drone and the payload by a perforated aluminum groundplane. Although the carbon fiber legs and struts are reflective across our frequency band, they retract shortly after takeoff and remain stowed behind the groundplane during flight. Range measurements of the assembled drone system (consisting of the drone, payload, groundplane, and transmitting antenna) indicate that the transmitting antenna has an omnidirectional shape, as expected, with full-width-half-max (FWHM) values of $60^{\circ}$ in the frequency ranges presented here. All measurements presented are within $3^{\circ}$ of the antenna pointing, and thus the effect of the transmission beam is expected to be negligible. As a result, we do not account for the transmission beam pattern here. The range measurements also indicated that the cross-pol from the transmitting antenna is typically -20dB or better and is featureless within the inner few degrees. We therefore expect that in the region where CHIME data is presented, the beam and polarization properties of the transmitting antenna can be assumed to be isotropic.
Our measurement is equivalent to a Ludwig-I polarization measurement instead of the more common Ludwig-III measurement in spherical coordinates [43]. This is a combination of two choices: first, the linear East-West flight path (described in more detail in Section 2.4) means that the drone did not maintain a constant radial distance from the center of CHIME. Second, the omnidirectional antenna is attached to the drone body without a gimbal mount and so the transmitted polarization axis is fixed to the drone body and thus the polarization angle remains parallel to CHIME during the course of the measurement. Although more complex flight paths and gimbal mounts[44] are an interesting direction for future measurements, the data presented here is restricted to the inner few degrees of the beam where these projection effects are <<1%. Most importantly, our measurements are in the near field without a transformation to the far-field, such that we do not present drone measurements with the interpretation that they are far-field CHIME beams.
@@ 2.3 CHIME Configuration for Drone Beammapping
[…]
The flashing drone calibration source is seen by all CHIME feeds, and the on- and off-times are differenced and then accumulated as an $N^{2}$ matrix of visibilities using the pulsar gating mode available from CHIME. To avoid smearing within an integration frame we must reduce the accumulation time from the standard 10 s down to 2 s, but this increased data rate of around 50 TB/hour can neither be offloaded from the CHIME X-engine nor can any useful length of data be stored within the on-site storage. To get around this we used the existing capabilities of the CHIME Science Data Processor (SDP) system to achieve a significant compression of the data[1].
For a source seen by all feeds, the resulting $N^{2}$ visibility matrix is effectively rank-1, meaning the majority of the signal power is captured in a single common mode. This common mode can be identified by performing a single-value decomposition (SVD 333numpy.linalg.svd documentation) on the matrix such that the common mode will be the first eigenmode of the matrix. Although any motion of the drone within an integration frame can generate additional modes, we found at 2 s cadence it is sufficient to record only a small number of the eigenvalues and vectors of the matrix to capture all of the information. Generating this eigendecomposition is computationally costly, and CHIME was designed to be able to do this once per 10 s frame for its source-based gain calibration. To fit within the computational footprint we only include data from a single cylinder of the instrument, which further reduces the output data rate. As we use an iterative algorithm for this calculation[1] this gives only a factor of 16 computational savings, rather than the factor of 64 savings we would expect for a full calculation.
The set of modified processing stages we perform on the $N^{2}$ matrix, per frequency, are:
  • Use the existing CHIME pulsar gating capability to take on-off differences of the calibration source signal on the full $N^{2}$ visibility matrix. This is possible as both CHIME’s clock and that of the noise source are locked to the GPS pulse-per-second (PPS) signal. This is then integrated to 2 s cadence.
@@
  • The existing SVD capability is used to calculate the largest 10 eigenvalues and their eigenvectors [1].
These eigenmodes, as well as the full visibility matrix for cylinder C, is then sent to the receiver node of the SDP system for the final processing, and writing of the following data products:
  • The full $N^{2}$ visibility matrix is written out for a set of 64 frequencies to allow us to verify that the eigendecomposition extracted all of the information.
  • For the full set of 1024 frequencies we write out the eigenvalues and eigenvectors to allow full sensitivity analysis of the data, as well as measurements of phase variation.
@@
The gated power for each feed can be reconstructed using the strongest 10 eigenmodes from a single cylinder. We found the ratio of the first eigenmode to the second mode was 100 or higher within the main beam, indicating that the majority ($99\%$) of the signal was the common mode response to the drone signal. As a result, the analysis in this paper uses the first eigenmode only, which sufficiently captures all of the drone signal while allowing for a lower noise floor than what is possible using the directly measured power contained in the autocorrelations.
@@ 2.4 Drone Flights
We performed ten flights over the CHIME telescope; in this paper we present results from the flight at the approximate declination of Virgo A, where independent beam measurements of CHIME from both holography data [33] and solar data [31] are available. Replicating the trajectory of Virgo A and the Sun through the CHIME primary beam required a 35 zenith angle drone flight relative to the array center. Due to altitude restrictions, we performed a low velocity flight at a conservative altitude to maximize the spatial resolution of the transit data. As shown in Figure 3, this flight was performed $\sim$220 m South from the center of the array, corresponding to distances between 180 - 250 m with respect to feed location along the focal line. Relative to the CHIME array center, the flight spanned -100 m to +100 m from East-to-West at an altitude of $\sim$307 m, in accordance with our flight ceiling limits from TransportCanada. The first flight was flown with a heading of $\sim 0^{\circ}$ [North] to measure co-polarizations corresponding to CHIME’s N (YY) polarization, and the second flight was flown with a heading of $\sim 90^{\circ}$ [East] to measure the orthogonal (XX) polarization. The typical distance from CHIME for these flights ranges from 350 m - 390 m. As a result, this flight plan corresponds to the near-field for all CHIME antennas (the CHIME far-field for a single feed is >>1 km). The time for each polarization pass was 3.4 mins.
@@ 3 DATA PROCESSING
Figure 3: The upper half of this figure shows coordinate conventions for the data from a drone flight at 35 degrees of zenith angle. One transit (upper left) across the beam of Cylinder C is shown true to scale in a local Cartesian coordinate system centered on the CHIME telescope. This flight trajectory is represented in three per-feed coordinate systems: Cartesian (upper right), Celestial (middle right) referenced to Local Meridian, and Orthographic (TelX/TelY)(lower right) using the color scale from Figure 2. The spherical coordinate system (bottom) with axes and angle definitions given to show the transformation from a spherical coordinate system to TelX (projection onto X-axis) and TelY (projection onto Y-axis).
@@ 3.1 Map Projections
During data processing, each CHIME correlation frame is associated with a concurrent drone position measurement from the RTK unit, which expresses the drone location in latitude, longitude, and height coordinates. A conversion to a local Cartesian reference frame [^4] PyGeodesy is performed using the latitude, longitude, and height of a known reference point (in this instance, the center of the CHIME array provided by surveying). In order to compare our measurements with other CHIME data sets, we must project our coordinates into a compatible reference frame, in particular, we choose TelX/TelY (described in more detail below) because it allowed us to compare to solar data results [31].
The drone flew a trajectory from East to West and then West to East, about 220 m from the center of CHIME, as shown in local Cartesian coordinates in Figure 3, with Cylinder C highlighted green. As shown in Figure 2, when the drone is directly on the boresight of CHIME cylinder, each feed will ‘see’ the drone at a slightly different angle up from the horizon (or down from zenith). This angle crucially depends on the altitude of the drone flight. The altitude must be measured relative to the phase center of the array (where incident rays are focused onto the feedline). We model the phase center of each feed by extending the incident ray to an image of the feedline placed beneath the primary reflector at a distance of 5 m (the focal length). This defines the image plane, also shown in Figure 2. The height of the CHIME cylinders above ground is $\sim$1 m, thus the image plane sits $\sim$4 m below the ground surface. As a result, if the drone reports a height of 307 m, we would expect to use 312 m (above the image plane) to find the correct angle. We will independently constrain the effective altitude in Section 3).
Once this drone altitude is established, we convert the positions surveyed during the drone flight from per-feed local Cartesian coordinates to TelX/TelY coordinates. The mapping of the drone flight path from local Cartesian to both celestial coordinates and TelX/TelY are shown in the right three panels of Figure 3, where the colorbar indicates feed number. The transformation from a spherical coordinate system to TelX/TelY is shown in the lower panel of Figure 3. In this coordinate system, the X-axis points East towards the horizon, and the Y-axis points North towards the horizon, and zenith (looking up) is defined along the Z-axis. TelX is the projection onto the X-axis, and represents the CHIME beam pattern parallel to the E-W axis; TelY is the projection onto the Y-axis, and represents the CHIME beam pattern parallel to the N-S axis. The resulting transformation is defined as:
$\mathrm{TelX}=\sin(\theta)\sin(\phi)$
$\mathrm{TelY}=\cos(\theta)$
Where $\theta$ is the angle from North and $\phi$ is the angle from Zenith measured in the XZ-plane. In this system, celestial sources will follow a curved path in TelX/TelY as can be seen in Figure 4. The choice of coordinate system was described in detail in [31], which presents solar data projected into the orthographic coordinate system (TelX, TelY). The advantage of this projection is that it orthogonalizes the features of the beam, and thus we use this projection throughout this paper.
@@ 3.2 Available Datasets
Comparisons are possible between three independent data sets (a) Drone measurements; (b) Solar measurements; and (c) Holography measurements. The overlap between these data sources can be seen in TelX/TelY coordinates in Figure 4 and given in Table 1. In the drone flights described above, because each feed will map to a different angle, the single East-West flight maps to a range of TelY values. As can be seen, the drone data overlaps the solar transit data and with one source measured via holography, Virgo A. All three data sets overlap at TelY $\sim-0.6$ which corresponds to a declination of $\sim 12^{\circ}$ in astronomical coordinates.
The drone data presented here are the first eigenmode (from the decomposition described in Section 2.3) for each feed on Cylinder C, and are measured in the near-field of CHIME. As described in[31], the solar data products are averages of cross-correlation visibilities of feeds within 10 m of each other, restricted to pairs of feeds within the same cylinder (intracylinder cross-correlations with baseline distances $\leq$10 m). These short baselines retain the solar signal, allowing an estimate for the CHIME primary beam averaged across all feeds and cylinders. The holography data[32] directly measures the far-field primary beam of each CHIME feed, and so feeds in Cylinder C can be selected and compared with the drone data. The holography data can also be processed with the same averaging methods used for the solar data for more direct comparisons across those two datasets. The parameter space spanned by these three data sets is given in Table 1.
Figure 4: Beam characterization measurements have been calculated from three unique data sets acquired in different portions of CHIME’s field of view. Coordinate axes are TelX and TelY, explained in more detail in Figure 3, which correspond to the projection to East and North, respectively. Solar beam measurements are presented in a continuous semicircular slice. Holographic measurements of galactic radio sources appear as pale yellow arcs. Measurements from the drone trajectory shown in Figure 3 are indicated by the square patch with solid yellow outline. Some additional drone flights (not analyzed here) are indicated by patches with dashed yellow outlines. The color scale of the solar and drone datasets is proportional to received power to illustrate the alignment of the main beam, but has not been appropriately calibrated for direct comparison. Comparison of the FWHM and Centroid measurements obtained from these data sets are shown in Figures 9 & 8.
Although the choice of TelX/TelY orthogonalizes the beam shape to align with the geometry of CHIME, sources at constant declination such as the holographic measurements of Virgo A and each day of solar data are curved in TelX/TelY (see Figure 4), which distorts sidelobe features. This distortion is appreciable outside $|\mathrm{TelX}|\gtrsim 0.25$ but negligible in the main beam where fits were performed. As a result, our determination of the centroid location and FWHM of the primary beam are not significantly distorted in the orthographic projection.
Data Set Solar Holography (transit) Drone
@@
Table 1: Survey coverage and parameter space of each data set. Virgo A for the holographic data set is in bold. More details are given in the text.